OpenStudy (anonymous):

WILL MEDAL CHECK.. Given that the first 4 terms of a geometric sequence are 2, x, 32, and 128, find the second term, x.3 ... ( I think the second term would be 4, )

OpenStudy (asnaseer):

Remember a geometric sequence is one where the ratio between successive terms in the sequence is the same. With this in mind, what is the ratio between the last two terms? i.e. what is: \(\frac{128}{32}\)

OpenStudy (anonymous):

is that 125 over 32?

OpenStudy (asnaseer):

128/32 - look at the last two terms of your sequence

OpenStudy (anonymous):

128/32 = 4

OpenStudy (asnaseer):

good - so now you know that you need to multiply by 4 to get to the next term

OpenStudy (asnaseer):

so what do you think x=?

OpenStudy (anonymous):

8

OpenStudy (acxbox22):

32/4=x

OpenStudy (asnaseer):

perfect! well done! :)

OpenStudy (anonymous):

another : Given that the first 5 terms of a geometric sequence are 3, x, 12, y, and 48, find, x and y. Assume both x and y are positive.

OpenStudy (triciaal):

I just answered that

OpenStudy (asnaseer):

<<<--- please ask each question separately in the question area on your left

OpenStudy (asnaseer):

this helps each question remain in focus

OpenStudy (triciaal):

go back to it and respond to my question